(-3, ___) is a solution of y = 4 - 3x
Now what I have tried is:
y = 4 - 3(-3)
y = 4 - -9
y = -5
So then I
y = 4 - 3(5)
y = 4 - 15
y = 9
I can't figure out this problem and I still don't understand how to come up with the two numbers needed

Answer by stanbon(75887)

(-3, ___) is a solution of y = 4 - 3x
Now what I have tried is:
y = 4 - 3(-3)
-------
y = 4 + 9
y = 13
----
Checking:
13 = 4 -3x
-3x = 9
x = 9/(-3) = -3
=====================
cheers,
Stan H.